# Manta: Enhancing Mamba for Few-Shot Action Recognition of Long Sub-Sequence

### v1

@@ Appendix C Acknowledgements
The authors would like to appreciate all participants of peer review and cloud servers provided by Paratera Ltd. Wenbo Huang sincerely thanks Bingxiao Shi (USTB) and all family members for the encouragement at an extremely difficult time. This work is supported by the National Key Research and Development Program for the 14th-Five-Year Plan of China 2023YFC3804104 in 2023YFC3804100; National Natural Science Foundation of China under Grants No.62072099, 62232004, 62373194, 62276063; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant BM2003201; Key Laboratory of Computer Network and Information Integration (MOE China) under Grant No.93K-9; the Fundamental Research Funds for the Central Universities; and JSPS KAKENHI Grant No.JP23K21676, JP24K02942, JP24K23849.

### v4

@@ Appendix C Acknowledgements
The authors would like to appreciate all participants of peer review and cloud servers provided by Paratera Ltd. Wenbo Huang sincerely thanks Bingxiao Shi (USTB), all family members, and TJUPT for the encouragement at an extremely difficult time. This work is supported by Frontier Technologies Research and Development Program of Jiangsu under Grant No. BF2024070; National Natural Science Foundation of China under Grants Nos. 62472094, 62072099, 62232004, 62373194, 62276063; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No. BM2003201; Key Laboratory of Computer Network and Information Integration (MOE, China) under Grant No. 93K-9; the Fundamental Research Funds for the Central Universities; the Excellent Ph.D Training Program (SEU); and JSPS KAKENHI Grant Nos. JP23K21676, JP24K02942, JP24K23849.
Codehttps://github.com/wenbohuang1002/Manta